OpenStudy (amenah8):

Find the volume of the solid generated when the region between the graphs f(x) = 2 sqrt(sinx) and y = 0 over the interval [0,pi] is revolved around the x-axis

satellite73 (satellite73):

that is just a picture of the region not the volume

satellite73 (satellite73):

itegrate \(\pi r^2\) (the area of the circle) from \(0\) to \(\pi\) in your case it should not be hard because \(r=f(x)=2\sqrt{\sin(x)}\)

satellite73 (satellite73):

*integrate

OpenStudy (amenah8):

the anti-derivative would be -4cosx+C

satellite73 (satellite73):

forget the C it is a definite integral

satellite73 (satellite73):

but yes, when you square and find the anti - derivative it is \(-4\cos(x)\)

satellite73 (satellite73):

oh don't forget the \(\pi\)

OpenStudy (amenah8):

so is the anser 4pi?

OpenStudy (amenah8):

because on my answer sheet they got 8pi

satellite73 (satellite73):

it is \(8\pi\)

satellite73 (satellite73):

\[-4\pi \cos(x)|^{\pi}_0\] \[=-4\pi(\cos(\pi)-\cos(0))\]

satellite73 (satellite73):

\[=-4\pi(-1-1)\] etc

OpenStudy (amenah8):

oh! i forgot to plug in 0

satellite73 (satellite73):

yeah you need that one too!

OpenStudy (amenah8):

thank you! that really helped